Hi, I am needing a little help with this algebra....
The base of a triangular sail is 18m shorter than it's height. The area is 140m^2. What is the base & height??
Thanks for any help, I really appreciate it!

Answer by nerdybill(7384)

The base of a triangular sail is 18m shorter than it's height. The area is 140m^2. What is the base & height??
.
For this problem, you're expected to know the area of any triangle is:
(1/2)bh
where
b is base
h is height or altitude
.
Let h = height
then
h-18 = base
.
140 = (1/2)(h-18)h
280 = h^2-18h
0 = h^2-18h-280
Factoring:
0 = (h+10)(h-28)
h = {28, -10}
Eliminate the negative answer:
h = 28 meters (height)
.
base:
h-18 = 28-18 = 10 meters (base)